Let 𝑦1 and 𝑦2 be linearly independent solutions of the differential equation 𝑦 β€²β€² + 𝑝(π‘₯)𝑦 β€² + π‘ž(π‘₯)𝑦 = 0, where functions 𝑝 and π‘ž are continuous on some interval 𝐼. (i) Prove that π‘Š(𝑦1, 𝑦2 )(π‘₯) = 𝐢𝑒 βˆ’ ∫ 𝑝(π‘₯)𝑑π‘₯ , where π‘Š is the Wronskian and 𝐢 ∈ ℝ is an arbitrary constant.



1
Expert's answer
2021-11-15T14:32:48-0500

Given that y1(x) and y2(x) are linear independent solutions of y''+p(x)y'+q(x)y=0 we get;

"y_1''+p(x)y_1'+q(x)y_1=0" .......(1)

"y_2''+p(x)y_2'+q(x)y_2=0" .......(2)

From (1) we get,

"y_1''=-(p(x)y_1'+y_1)" ......(3)

And from (2),

"y_2''=-(p(x)y_2'+y_2)" ........(4)

We know that the Wronskian

"W(y_1,y_2)=\begin{vmatrix}
 y_1 & y_2 \\
 y_1' & y_2'
\end{vmatrix}=y_1y_2'-y_2y_1'"

"W'=(y_1y_2'-y_2y_1')'"

"W'=(y_1y_2')'-(y_2y_1')'"

We know "(uv)'=u\cdot v'+v\cdot u'"

"\therefore W'=(y_1\cdot y_2''+y_2'\cdot y_1')-(y_2\cdot y_1''+y_1'\cdot y_2')"

"W'=y_1\cdot y_2''+y_2'\cdot y_1'-y_2\cdot y_1''-y_1'\cdot y_2'"

"W'=y_1[-(p(x)y_2'+y_2)]-y_2[-(p(x)y_1'+y_1)]"

Values from (1) and (2)

"W'=-p(x)y_1y_2'-y_1y_2+p(x)y_2y_1'+y_1y_2"

"W'=-p(x)y_1y_2'+p(x)y_2y_1"

"W'=-p(x)[y_1y_2'-y_2y_1']"

"W'=-p(x)W"

Which is an equation of the form z'=-pz whose solution is of the form "z=ce^{-\int Pdx}"

From "W'=-pW" we get;

"W(y_1,y_2)=ce^{-\int Pdx}"